Shawnee Motors Inc. assembles and sells MP3 players. The company began operations on August 1 and operated at 100% of capacity during the first month. The following data summarize the results for August: Sales (14,500 units) Production costs (19,000 units): $2,175,000 Direct materials Direct labor Variable factory overhead Fixed factory overhead $1,054,500 505,400 252,700 169,100 1,981,700 Selling and administrative expenses Variable selling and administrative expenses $307,200 Fixed selling and administrative expenses 118,900 426,100 If required, round interim per-unit calculations to the nearest cent. a. Prepare an income statement according to the absorption costing concept. Shawnee Motors Inc. Absorption Costing Income Statement For the Month Ended August 31 Sales Cost of goods soldGross profit Selling and administrative expenses Income from operations eedback b. Prepare an income statement according to the variable costing concept. Shawnee Motors Inc. Variable Costing Income Statement For the Month Ended August 31 Sales Variable cost of goods sold Manufacturing margin Variable selling and administrative expenses Contribution margin Fixed costs: Fixed factory overhead Fixed selling and administrative expenses Total fixed costs Income from operations
Answer #1

Absorption costing Income Statement:

For absorption costing all the production costs including the fixed costs are considered. Therefore, for preparing absorption cost income statement, it is required to find the cost per unit, which is calculated as follows:

Direct Material - 1054500/19000 = 55.5

Direct Labor = 505400/19000 = 26.6

Variable factory Overhead = 252700/19000 = 13.3

Fixed Factory Overhead = 169100/19000 = 8.9

Cost of Goods sold per unit = $104.30

Variable Selling and administrative expenses per unit = 307200/19000 = 16.17

Total variable selling and administrative expenditure - 16.17*14,500 units + 118,900 = 353,365

Absorption Costing Income statement:

Sales (14,500 units) = 2,175,000

Less:Cost of Goods Sold (14500*104.30) = 1,512,350

Gross Profit = 662,650

Less : total selling and administrative expenses=   353,365

Operating Income = 309,285

Variable Costing Income Statement:

Sales(14,500 units) = 2,175,000

Less: Variable Cost (14,500 units)

Direct Material = 804.750

Direct Labor = 385,700

Variable Factory Overhead = 192,850

Variable selling and administrative Expenses = 234,465

Contribution = 557,235

Less : Fixed Costs

Fixed Factory Overhead = 307,200

Fixed selling and admin expenses = 118,900

Operating Income = 131,135